A chair of weight 75.0 N lies atop a horizontal floor; the floor is not frictionless. You push on the chair with a force of F = 43.0 N directed at an angle of 40.0 degrees below the horizontal and the chair slides along the floor.

Using Newton's laws, calculate n, the magnitude of the normal force that the floor exerts on the chair.

Express your answer in newtons.
Answer #1

Magnitude of the normal force

Chair is stationary in the y direction

F_y = 0

n - w - F*sin 40° = 0

n = w + F*sin40°

= 75 + (43)sin40°  

n = 102.6 N